OpenStudy (bbb911):

This is Geometry: Find the scale factor of ^DEF to ^XYZ, DE= 10, XZ = 20, and XY = 45.

OpenStudy (anonymous):

45/10=4.5

OpenStudy (bbb911):

:)

OpenStudy (bbb911):

Here is a mutliple choice for the question A.1:2 B. 1:4 C. 2:9 D. 2:11

OpenStudy (anonymous):

2/9

OpenStudy (anonymous):

C

OpenStudy (anonymous):

4.5=9/2=9:2 so C

OpenStudy (bbb911):

thanks you both